In an adult, the aorta's radius is normally ~1.5cm, and blood moves through it at an average speed of 30m/s. If a typical capillary has a radius of 5×10-6m and blood passes through them with velocity of 0.1cm/s approximately how many capillaries are in the blood?


1
Expert's answer
2021-11-23T10:32:28-0500

The volumetric flows that pass through the aorta and capillaries are equal:


"A_1v_1=A_2v_2\equiv\text{const}.\\\space\\
A_a=\pi R^2,A_c=\pi r^2,\\\space\\
A_av_a=NA_cv_c,\\\space\\
N=\frac{A_av_a}{A_cv_c}=\frac{R^2v_a}{r^2v_c},\\\space\\
N=\frac{0.015^2·30}{(5·10^{-6})^2·0.001}=2.7·10^{11}."